OpenStudy (anonymous):

If object moves on line with velocity v(t), where v(0) = 1 and v(1) = 2. And there is a constant k such that a = k*(t^3/v^2). Find velocity of object at time t=2

OpenStudy (anonymous):

v = v(t)m/s, a = a(t) m/s^2 v(0) = 1 v(1) = 2 \[a = k \frac{ t ^{3} }{ v^{2} }\] Find v(2)

OpenStudy (anonymous):

I think I may have to use this equation: \[v(t) = \int\limits_{}^{}a(t)dt + k\]